A hurricane wind blows across a 7.50 m × 11.2 m flat roof at a speed of 190 km/hr.

a) Is the air pressure above the roof higher or lower than the pressure inside the house? Explain.
b) What is the pressure difference?
c) How much force is exerted on the roof?
d) If the roof cannot withstand this much force, will it “blow in” or “blow out”?
Answer #1
a.
the pressure is lower on the roof as a consequence of bernoulli's theorem
it says p + 0.5 d*v^2 = constant
more the speed less the pressure
implies inside is more pressure
on the roof the speed of wind is more so less pressure

speed = 190*5/18 = 52.78 m/s^2
b. pressure difference = 1/2 density * v^2 = 1.25 * 0.5 * 52.78^2 = 1742 N/m^2

c. force = area * pressure = 1742* 7.5*11.2 = 146328 N

d. the roof will blow out because pressure is more in the inside